# Graph Retrieval Augmented Trustworthiness Reasoning

### v1

@@ Introduction
In multiplayer games with incomplete information, reasoning is crucial for assessing the trustworthiness of players who may conceal their intentions based on their actions, dialogue, and other observable information. Autonomous agents must analyze and evaluate the reliability of available information sources to determine player trust and cooperation (Fig. 1). Currently, key methods supporting such reasoning include symbolic reasoning (nye2021improving ; coherence and consistency in neural sequence models with dual system 2021) and evidential theory (geometry problem solving with formal language and symbolic reasoning 2021), Bayesian reasoning (hierarchy hesitant fuzzy linguistic entropy-based TODIM approach using evidential theory 2020; probability to consilience: How explanatory values implement Bayesian reasoning 2021; implementations of Bayesian inference 2021), and reinforcement learning (RL) (a bayesian multi-hop reasoning framework for knowledge graph reasoning 2021; like human: Hierarchical reinforcement learning for knowledge graph reasoning 2020; attention-based deep reinforcement learning framework for knowledge graph reasoning 2021).
@@
Retrieval augmented generation (RAG)(aware knowledge graph reasoning based on deep reinforcement learning 2024; for Large Language Models: A Survey 2024) is an alternative paradigm for addressing LLM limitations by incorporating an external retrieval mechanism. In the RAG pipeline, a retriever first indexes and retrieves relevant data chunks, which are then used in conjunction with an input query to enhance the generation process. Optimization techniques such as step-by-step retrieval using a Chain of Thought (CoT) (for AI-Generated Content: A Survey 2024) decomposition of the query, fine-tuning embedding models to capture semantically similar content in vector space, and combining various retrieval methods to improve the robustness of the retriever. This approach can help mitigate issues of knowledge obsolescence and hallucinations by integrating up-to-date and contextually relevant information.
@@
In incomplete information games, players typically need to reason to get more potential information based on the observations in order to make better decisions. Especially in multiplayer games, players need to observe and analyze other players’ behavior and historical data in real time. Typically, there is a large amount of data to be processed, and some of it contains misleading information. Hence, the reasoning ability of agents is of crucial importance (augmented thoughts elicit context-aware reasoning in long-horizon generation 2024; for Large Language Models in the Game Werewolf 2024; based Agents for Large-Scale Decision-Making: An Actor-Critic Approach 2024; playing Adversarial Language Game Enhances LLM Reasoning 2024; for Evaluating Sequential Decision-Making Capability of Large Language Models 2024).
In traditional game theoretic algorithms, the game is usually played through Bayesian method (of LLM Agents 2020) and evolutionary game theory (games: Games with incomplete information 2015). In addition, there are machine learning methods such as Monte Carlo tree search (combination from an evolutionary game theory perspective 2012) and RL (set monte carlo tree search 2016). With the introduction of deep networks, RL methods have been widely used in multiplayer games due to their excellent inference ability.
While RL offer significant advantages in reasoning, it require a large amount of domain-specific training data and their inference capabilities are limited to that domain, making them unsuitable for a wide range of reasoning tasks. LLM-based reasoning provide new ideas for completing reasoning tasks due to their extensive knowledge and powerful language capabilities. Xu et.al. (reinforcement learning from self-play in imperfect-information games 2023) proposed a framework combining LLMs and RL is proposed to develop strategic language agents that can overcome inherent biases and perform well in complex decision-making tasks by allowing RL strategies to choose the optimal decision from a diverse set of action candidates generated by LLMs.
@@
However, not all directly retrieved document chunks are highly relevant to the query. Reranking algorithms (agents with reinforcement learning for strategic play in the werewolf game 2023) address this by reordering the retrieved chunks, allowing the generator to make better use of the most relevant information and generate responses that are more accurate and pertinent to the query.
The Graph RAG method (good at search? investigating large language models as re-ranking agents 2024) further improves this process by constructing a graph model from the retrieved document chunks, representing entities and relationships as nodes and directed edges. Subsequently, knowledge and relationships related to the query are retrieved based on this graph, thus assisting the LLM in generating a more relevant response.
@@
In this section, we choose Xu’s algorithm (local to global: A graph rag approach to query-focused summarization 2024) as the baseline LLM. We validate our approach through experiments on the multiplayer game ”Werewolf”, comparing GRATR with baseline LLM and LLMs enhanced with Native RAG and Rerank RAG in terms of win rates, performance, and reasoning ability. The source code of the proposed GRATR is available at https://github.com/EvoNexusX/Graph-Retrieval-Augmented-Trustworthiness-Reasoning .
@@
We implemented our GRATR method using the classic multi-player game ”Werewolf”. We used the framework of the Werewolf environment implemented by Ref (local to global: A graph rag approach to query-focused summarization 2024). We set the number of players to 8 in the Werewolf game, with three leaders: the witch, the guard, and the seer, three werewolves, and two villagers. The history message window size $K$ is set to 15. We use the gpt-3.5-turbo model as the backend LLM, and its temperature is set to 0.3.
@@ Appendix 3. Summary and Augment Reasoning
Appendix C Checklist
  1. 1.
    This paper:
    1. (a)
      Includes a conceptual outline and/or pseudocode description of AI methods introduced yes
    2. (b)
      Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results yes
    3. (c)
      Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper yes
  2. 2.
    Does this paper make theoretical contributions? no
  3. 3.
    Does this paper rely on one or more datasets? no
  4. 4.
    Does this paper include computational experiments? yes
    1. (a)
      Any code required for pre-processing data is included in the appendix. yes
    2. (b)
      All source code required for conducting and analyzing the experiments is included in a code appendix. no
    3. (c)
      All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. yes
    4. (d)
      All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from yes
    5. (e)
      If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. NA
    6. (f)
      This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. no
    7. (g)
      This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. yes
    8. (h)
      This paper states the number of algorithm runs used to compute each reported result. yes
    9. (i)
      Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. no
    10. (j)
      The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). no This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. yes
    11. (k)
      This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. no

### v2

@@ Introduction
In multiplayer games with incomplete information, reasoning is crucial for assessing the trustworthiness of players who may conceal their intentions based on their actions, dialogue, and other observable information. Autonomous agents must analyze and evaluate the reliability of available information sources to determine player trust and cooperation (Fig. 1). Currently, key methods supporting such reasoning include symbolic reasoning (Nye et al. 2021; Lu et al. 2021) and evidential theory (Liu et al. 2021), Bayesian reasoning (Wojtowicz and DeDeo 2020; Sohn and Narain 2021; Wan and Du 2021), and reinforcement learning (RL) (Wan et al. 2021; Wang, Hao, and Cao 2020; Tiwari, Zhu, and Pandey 2021).
@@
Retrieval augmented generation (RAG)(Gao et al. 2024; Zhao et al. 2024) is an alternative paradigm for addressing LLM limitations by incorporating an external retrieval mechanism. In the RAG pipeline, a retriever first indexes and retrieves relevant data chunks, which are then used in conjunction with an input query to enhance the generation process. Optimization techniques such as step-by-step retrieval using a Chain of Thought (CoT) (Wang et al. 2024) decomposition of the query, fine-tuning embedding models to capture semantically similar content in vector space, and combining various retrieval methods to improve the robustness of the retriever. This approach can help mitigate issues of knowledge obsolescence and hallucinations by integrating up-to-date and contextually relevant information.
@@
In incomplete information games, players typically need to reason to get more potential information based on the observations in order to make better decisions. Especially in multiplayer games, players need to observe and analyze other players’ behavior and historical data in real time. Typically, there is a large amount of data to be processed, and some of it contains misleading information. Hence, the reasoning ability of agents is of crucial importance (Wu et al. 2024; Zhang et al. 2024; Cheng et al. 2024; Qin et al. 2024; Costarelli et al. 2024).
In traditional game theoretic algorithms, the game is usually played through Bayesian method (Zamir 2020) and evolutionary game theory (Deng et al. 2015). In addition, there are machine learning methods such as Monte Carlo tree search (Cowling, Powley, and Whitehouse 2012) and RL (Heinrich and Silver 2016). With the introduction of deep networks, RL methods have been widely used in multiplayer games due to their excellent inference ability.
While RL offer significant advantages in reasoning, it require a large amount of domain-specific training data and their inference capabilities are limited to that domain, making them unsuitable for a wide range of reasoning tasks. LLM-based reasoning provide new ideas for completing reasoning tasks due to their extensive knowledge and powerful language capabilities. Xu et.al. (Xu et al. 2023) proposed a framework combining LLMs and RL is proposed to develop strategic language agents that can overcome inherent biases and perform well in complex decision-making tasks by allowing RL strategies to choose the optimal decision from a diverse set of action candidates generated by LLMs.
@@
However, not all directly retrieved document chunks are highly relevant to the query. Reranking algorithms (Sun et al. 2023) address this by reordering the retrieved chunks, allowing the generator to make better use of the most relevant information and generate responses that are more accurate and pertinent to the query.
The Graph RAG method (Edge et al. 2024) further improves this process by constructing a graph model from the retrieved document chunks, representing entities and relationships as nodes and directed edges. Subsequently, knowledge and relationships related to the query are retrieved based on this graph, thus assisting the LLM in generating a more relevant response.
@@
In this section, we choose Xu’s algorithm (Xu et al. 2024) as the baseline LLM. We validate our approach through experiments on the multiplayer game ”Werewolf”, comparing GRATR with baseline LLM and LLMs enhanced with Native RAG and Rerank RAG in terms of win rates, performance, and reasoning ability. The source code of the proposed GRATR is available at https://github.com/EvoNexusX/Graph-Retrieval-Augmented-Trustworthiness-Reasoning .
@@
We implemented our GRATR method using the classic multi-player game ”Werewolf”. We used the framework of the Werewolf environment implemented by Ref (Xu et al. 2024). We set the number of players to 8 in the Werewolf game, with three leaders: the witch, the guard, and the seer, three werewolves, and two villagers. The history message window size $K$ is set to 15. We use the gpt-3.5-turbo model as the backend LLM, and its temperature is set to 0.3.